Consider using a z test to test H0: p = 0.5.

Determine the P-value in each of the following situations. (Round your answers to four decimal places.)

(a) Ha: p > 0.5, z = 1.45

(b) Ha: p < 0.5, z = −2.76

(c) Ha: p ≠ 0.5, z = −2.76

(d) Ha: p < 0.5, z = 0.21
Answer #1

Solution :

Given that,

The null and alternative hypothesis is

H0 : p =0.5

a) Ha: p > 0.5, z = 1.45

This is the right tailed test

Using standard normal table,  

P(Z > 1.45)

= 1 - P( Z < 1.45)

= 1 - 0.9265

p-value = 0.0735

b) Ha: p < 0.5, z = −2.76

This is the left tailed test

Using standard normal table,

P( Z < -2.76)

p-value = 0.0029

c) Ha: p ≠ 0.5, z = −2.76

This is the two tailed test

Using standard normal table,

P( Z < -2.76)

= 0.0029

p-value = 2 * P( Z < -2.76)

p-value = 2 * 0.0029

p-value = 0.0058

d) Ha: p < 0.5, z = 0.21

This is the left tailed test

P(Z > 0.21)

= 1 - P( Z < 0.21)

= 1 - 0.5831

p-value = 0.4169